Mrs. Tom has a son and a duaghter. She is twice as old as her son and her son is twice as old as her daughter. The total of thier ages is 91 years. How old is her son?
Answer by mananth(16946)

daughter age = x
son's age = 2x
mother's age = 2*2x=4x
x+2x+4x=91
7x=91
x=13
daughter = 13
son =2x=> 26
mother = 4x=> 52